# Discriminating image representations with principal distortions

### v1

@@ Acknowledgements
6 Reproducibility Statement
The code used to generate principal distortions and details of loading the models is uploaded as a .zip file for the submission, and will be made available via a public GitHub repository upon publication. We aim for the distortion generation to be easy for others to use to probe new models. The mathematical derivation of the algorithm is provided in Appx. B. We have included details in Appx. C with the parameters of the early visual models, and Appx. D details where checkpoints were obtained for the tested deep neural networks.
@@ Appendix A Relation to the Fisher-Rao metric
$\frac{1}{K}(\log\lambda_{1}-\log\lambda_{K})^{2}\leq\gamma^{2}({\bm{A}},{\bm{B}})\leq\frac{K-1}{2}(\log\lambda_{1}-\log\lambda_{K})^{2}.$
When ${\bm{A}}={\bm{I}}_{A}$ and ${\bm{B}}={\bm{I}}_{B}$, then $d_{A}({\bm{\epsilon}})=\sqrt{{\bm{\epsilon}}^{\top}{\bm{A}}{\bm{\epsilon}}}$ and $d_{B}({\bm{\epsilon}})=\sqrt{{\bm{\epsilon}}^{\top}{\bm{B}}{\bm{\epsilon}}}$. If ${\bm{\epsilon}}_{1}$ and ${\bm{\epsilon}}_{K}$ denote the extremal generalized eigenvectors associated with $\lambda_{1}$ and $\lambda_{K}$, respectively, then
@@
$\frac{2}{\sqrt{K}}m_{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{K}}({\bm{I}}_{A},{\bm{I}}_{B})\leq\gamma({\bm{A}},{\bm{B}})\leq\sqrt{2(K-1)}m_{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{K}}({\bm{I}}_{A},{\bm{I}}_{B}).$
@@
Equation 5 suggests a natural extension for defining a metric between positive definite matrices using $M>2$ distortions. Specifically, it suggests choosing the distortions to maximize the variance across the log ratios of the sensitivities. To this end, we can define the squared distance between the local geometries to be the variance (across distortions) of the log ratio of the the sensitivities:
$m_{{\bm{\epsilon}}_{1},\dots,{\bm{\epsilon}}_{M}}^{2}({\bm{A}},{\bm{B}})=M\text{Var}\left(\left\{\log\frac{d_{A}({\bm{\epsilon}}_{m})}{d_{B}({\bm{\epsilon}}_{m})}\right\}\right).$
A set of $M$ principal distortions can then be chosen to maximize the variance across models under this metric.
@@ Appendix B Computing the top two optimal distortions
$\operatorname*{arg\,max}_{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}}L({\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}),$ $L({\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}):=\sum_{n=1}^{N}\left\{\log\frac{d_{n}({\bm{\epsilon}}_{1})}{d_{n}({\bm{\epsilon}}_{2})}-\frac{1}{N}\sum_{m=1}^{N}\log\frac{d_{m}({\bm{\epsilon}}_{1})}{d_{m}({\bm{\epsilon}}_{2})}\right\}^{2}.$
Differentiating $L$ with respect to ${\bm{\epsilon}}_{1}$ yields
$\nabla_{{\bm{\epsilon}}_{1}}L({\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2})$ $=2\sum_{n=1}^{N}\left\{\log\frac{d_{n}({\bm{\epsilon}}_{1})}{d_{n}({\bm{\epsilon}}_{2})}-\frac{1}{N}\sum_{m=1}^{N}\log\frac{d_{m}({\bm{\epsilon}}_{1})}{d_{m}({\bm{\epsilon}}_{2})}\right\}\left\{\frac{{\bm{I}}_{n}({\mathbf{s}}){\bm{\epsilon}}_{1}}{d_{n}^{2}({\bm{\epsilon}}_{1})}-\frac{1}{N}\sum_{m=1}^{N}\frac{{\bm{I}}_{m}({\mathbf{s}}){\bm{\epsilon}}_{1}}{d_{m}^{2}({\bm{\epsilon}}_{1})}\right\}$
$=\sum_{n=1}^{N}\left\{\log\frac{d_{n}^{2}({\bm{\epsilon}}_{1})}{d_{n}^{2}({\bm{\epsilon}}_{2})}-\frac{1}{N}\sum_{m=1}^{N}\log\frac{d_{m}^{2}({\bm{\epsilon}}_{1})}{d_{m}^{2}({\bm{\epsilon}}_{2})}\right\}\left\{\frac{{\bm{I}}_{n}({\mathbf{s}}){\bm{\epsilon}}_{1}}{d_{n}^{2}({\bm{\epsilon}}_{1})}-\frac{1}{N}\sum_{m=1}^{N}\frac{{\bm{I}}_{m}({\mathbf{s}}){\bm{\epsilon}}_{1}}{d_{m}^{2}({\bm{\epsilon}}_{1})}\right\},$
@@
$\nabla_{{\bm{\epsilon}}}\log d({\bm{\epsilon}})$ $=\frac{1}{2}\nabla_{{\bm{\epsilon}}}\log({\bm{\epsilon}}^{\top}{\bm{I}}{\bm{\epsilon}})=\frac{{\bm{I}}{\bm{\epsilon}}}{{\bm{\epsilon}}^{\top}{\bm{I}}{\bm{\epsilon}}}=\frac{{\bm{I}}{\bm{\epsilon}}}{d^{2}({\bm{\epsilon}})}.$
Similarly, differentiating $L$ with respect to ${\bm{\epsilon}}_{2}$ yields:
$\nabla_{{\bm{\epsilon}}_{2}}L({\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2})$ $=-\sum_{n=1}^{N}\left\{\log\frac{d_{n}^{2}({\bm{\epsilon}}_{1})}{d_{n}^{2}({\bm{\epsilon}}_{2})}-\frac{1}{N}\sum_{m=1}^{N}\log\frac{d_{m}^{2}({\bm{\epsilon}}_{1})}{d_{m}^{2}({\bm{\epsilon}}_{2})}\right\}\left\{\frac{{\bm{I}}_{n}({\mathbf{s}}){\bm{\epsilon}}_{2}}{d_{n}^{2}({\bm{\epsilon}}_{2})}-\frac{1}{N}\sum_{m=1}^{N}\frac{{\bm{I}}_{m}({\mathbf{s}}){\bm{\epsilon}}_{2}}{d_{m}^{2}({\bm{\epsilon}}_{2})}\right\}.$
@@
2:  Initialize: distortions ${\bm{\epsilon}}_{1},{\bm{\epsilon}}_{2}\in\mathbb{R}^{D}$
@@
5:        ${\bm{v}}_{1}(n)\leftarrow{\bm{I}}_{n}{\bm{\epsilon}}_{1}$
6:        ${\bm{v}}_{2}(n)\leftarrow{\bm{I}}_{n}{\bm{\epsilon}}_{2}$
7:        $d_{1}^{2}(n)\leftarrow\langle{\bm{\epsilon}}_{1},{\bm{v}}_{1}(n)\rangle$
8:        $d_{2}^{2}(n)\leftarrow\langle{\bm{\epsilon}}_{2},{\bm{v}}_{2}(n)\rangle$
@@
16:     $\Delta{\bm{\epsilon}}_{1}\leftarrow\sum_{n=1}^{N}\left[r(n)-\bar{r}\right]\left[{\bm{u}}_{1}(n)-\bar{\bm{u}}_{1}\right]$
17:     $\Delta{\bm{\epsilon}}_{2}\leftarrow-\sum_{n=1}^{N}\left[r(n)-\bar{r}\right]\left[{\bm{u}}_{2}(n)-\bar{\bm{u}}_{2}\right]$
18:     ${\bm{\epsilon}}_{1}\leftarrow{\bm{\epsilon}}_{1}+\eta\Delta{\bm{\epsilon}}_{1}$
19:     ${\bm{\epsilon}}_{2}\leftarrow{\bm{\epsilon}}_{2}+\eta\Delta{\bm{\epsilon}}_{2}$
20:     ${\bm{\epsilon}}_{1}\leftarrow\alpha{\bm{\epsilon}}_{1}/\|{\bm{\epsilon}}_{1}\|$
21:     ${\bm{\epsilon}}_{2}\leftarrow\alpha{\bm{\epsilon}}_{2}/\|{\bm{\epsilon}}_{2}\|$
@@ Appendix D Methods: Deep Neural Network Experiments
The analyzed DNNs were obtained from the model loading code and checkpoints available at https://github.com/jenellefeather/model_metamers_pytorch which were used in Feather et al. (2023), and allowed for easy loading and selection of the intermediate layer stages for many models that had previously been proposed as models of human visual perception. The checkpoints for the standard ResNet50 and AlexNet models were obtained from the public pytorch checkpoints ((Ansel et al., 2024)). The Stylized Image Net AlexNet (alexnet_trained_on_SIN) and ResNet50 (resnet50_trained_on_SIN) were obtained from https://github.com/rgeirhos/texture-vs-shape associated with Geirhos et al. (2019). The checkpoint for the ResNet50 $\ell_{2}(\epsilon_{\text{train}}=3.0)$ adversarially trained model was obtained from Engstrom et al. (2019), and the checkpoint for the Alexnet $\ell_{2}(\epsilon_{\text{train}}=3.0)$ adversarially trained model was obtained from Feather et al. (2023). For all experiments, we only included intermediate layers before the final classification stage in the principal distortion analysis, and the subset of layers followed those chosen in Feather et al. (2023). Specifically, for the AlexNet models we included layers relu0, relu1, relu2, relu3, relu4, fc0_relu, and fc1_relu in each set of anlayses. For the ResNet50 models we included conv1_relu1, layer1, layer2, layer3, layer4, and avgpool in each set of analyses.
D.1 Principal distortion optimization
@@
For the deep neural network experiments, we use a subset of 100 ImageNet images where each image was chosen from a unique class (randomly chosen from the set of images at https://github.com/EliSchwartz/imagenet-sample-images). For each image and each comparison, we ran the gradient descent procedure for principal distortion optimization for 2500 iterations, using an exponentially decaying learning rate that started at 10.0 and decayed to 0.001 by the final step.
@@
We used a target distortion size of $\alpha=0.1$ and at each step of the optimization, we also scaled ${\bm{\epsilon}}$ so that the image ${\mathbf{s}}+1000{\bm{\epsilon}}$ would not be clipped when the RGB value was represented between 0–1; that is, we scaled ${\bm{\epsilon}}$ so that $0\leq s[i]+1000\epsilon[i]\leq 1$ for each value $i$ in the image. This constraint could potentially bias the perturbations to be more spread out across the image (because all of the amplitude for the perturbation cannot be focused at a small set of pixels). However, removing this constraint did not lead to quantitatively different results (Supplementary Fig. 10), but the inclusion allows for more viable comparisons with human perception (as the perturbation can be scaled while maintaining valid values in the image gamut).
D.2 Supplemental Figures

### v2

@@ Acknowledgments
Reproducibility Statement
The code used to generate principal distortions, details of loading the models, and example principal distortion generation code is available at
https://github.com/LabForComputationalVision/principal-distortions
We aim for the distortion generation to be easy for others to use to probe new models. The mathematical derivation of the algorithm is provided in Appx. B. We have included details in Appx. C with the parameters of the early visual models, and Appx. D details where checkpoints were obtained for the tested deep neural networks.
@@ Appendix A Relation to the Fisher-Rao metric
$\frac{1}{K}(\log\lambda_{1}-\log\lambda_{K})^{2}\leq\gamma^{2}({\bm{A}},{\bm{B}})\leq\frac{K-1}{2}(\log\lambda_{1}-\log\lambda_{K})^{2},$
with equality holding when $K=2$. When ${\bm{A}}={\bm{I}}_{A}$ and ${\bm{B}}={\bm{I}}_{B}$, then $d_{A}({\bm{\epsilon}})=\sqrt{{\bm{\epsilon}}^{\top}{\bm{A}}{\bm{\epsilon}}}$ and $d_{B}({\bm{\epsilon}})=\sqrt{{\bm{\epsilon}}^{\top}{\bm{B}}{\bm{\epsilon}}}$. If ${\bm{\epsilon}}_{1}$ and ${\bm{\epsilon}}_{K}$ denote the extremal generalized eigenvectors associated with $\lambda_{1}$ and $\lambda_{K}$, respectively, then
@@
$\frac{2}{\sqrt{K}}m_{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{K}}({\bm{I}}_{A},{\bm{I}}_{B})\leq\gamma({\bm{A}},{\bm{B}})\leq\sqrt{2(K-1)}m_{{\bm{\epsilon}}_{1},{\bm{\epsilon}}_{K}}({\bm{I}}_{A},{\bm{I}}_{B}).$ (6)
@@
Equation 5 suggests a natural extension for defining a metric between positive definite matrices using $J>2$ distortions. Specifically, it suggests choosing the distortions to maximize the variance across the log ratios of the sensitivities. To this end, we can define the squared distance between the local geometries to be the variance (across distortions) of the log ratio of the the sensitivities:
$m_{{\mathbf{u}}_{1},\dots,{\mathbf{u}}_{J}}^{2}({\bm{A}},{\bm{B}})=M\text{Var}\left(\left\{\log\frac{d_{A}({\mathbf{u}}_{j})}{d_{B}({\mathbf{u}}_{j})}\right\}\right).$
A set of $J$ principal distortions can then be chosen to maximize the variance across models under this metric.
@@ Appendix B Computing the top two optimal distortions
$\operatorname*{arg\,max}_{{\mathbf{u}},{\mathbf{v}}}L({\mathbf{u}},{\mathbf{v}}),$ $L({\mathbf{u}},{\mathbf{v}}):=\sum_{n=1}^{N}\left\{\log\frac{d_{n}({\mathbf{u}})}{d_{n}({\mathbf{v}})}-\frac{1}{N}\sum_{m=1}^{N}\log\frac{d_{m}({\mathbf{u}})}{d_{m}({\mathbf{v}})}\right\}^{2}.$
Differentiating $L$ with respect to ${\mathbf{u}}$ yields
$\nabla_{{\mathbf{u}}}L({\mathbf{u}},{\mathbf{v}})$ $=2\sum_{n=1}^{N}\left\{\log\frac{d_{n}({\mathbf{u}})}{d_{n}({\mathbf{v}})}-\frac{1}{N}\sum_{m=1}^{N}\log\frac{d_{m}({\mathbf{u}})}{d_{m}({\mathbf{v}})}\right\}\left\{\frac{{\bm{I}}_{n}({\mathbf{s}}){\mathbf{u}}}{d_{n}^{2}({\mathbf{u}})}-\frac{1}{N}\sum_{m=1}^{N}\frac{{\bm{I}}_{m}({\mathbf{s}}){\mathbf{u}}}{d_{m}^{2}({\mathbf{u}})}\right\}$
$=\sum_{n=1}^{N}\left\{\log\frac{d_{n}^{2}({\mathbf{u}})}{d_{n}^{2}({\mathbf{v}})}-\frac{1}{N}\sum_{m=1}^{N}\log\frac{d_{m}^{2}({\mathbf{u}})}{d_{m}^{2}({\mathbf{v}})}\right\}\left\{\frac{{\bm{I}}_{n}({\mathbf{s}}){\mathbf{u}}}{d_{n}^{2}({\mathbf{u}})}-\frac{1}{N}\sum_{m=1}^{N}\frac{{\bm{I}}_{m}({\mathbf{s}}){\mathbf{u}}}{d_{m}^{2}({\mathbf{u}})}\right\},$
@@
$\nabla_{{\mathbf{u}}}\log d({\mathbf{u}})$ $=\frac{1}{2}\nabla_{{\mathbf{u}}}\log({\mathbf{u}}^{\top}{\bm{I}}{\mathbf{u}})=\frac{{\bm{I}}{\mathbf{u}}}{{\mathbf{u}}^{\top}{\bm{I}}{\mathbf{u}}}=\frac{{\bm{I}}{\mathbf{u}}}{d^{2}({\mathbf{u}})}.$
Similarly, differentiating $L$ with respect to ${\mathbf{v}}$ yields:
$\nabla_{{\mathbf{v}}}L({\mathbf{u}},{\mathbf{v}})$ $=-\sum_{n=1}^{N}\left\{\log\frac{d_{n}^{2}({\mathbf{u}})}{d_{n}^{2}({\mathbf{v}})}-\frac{1}{N}\sum_{m=1}^{N}\log\frac{d_{m}^{2}({\mathbf{u}})}{d_{m}^{2}({\mathbf{v}})}\right\}\left\{\frac{{\bm{I}}_{n}({\mathbf{s}}){\mathbf{v}}}{d_{n}^{2}({\mathbf{v}})}-\frac{1}{N}\sum_{m=1}^{N}\frac{{\bm{I}}_{m}({\mathbf{s}}){\mathbf{v}}}{d_{m}^{2}({\mathbf{v}})}\right\}.$
@@
2:  Initialize: distortions ${\mathbf{u}},{\mathbf{v}}\in\mathbb{R}^{D}$
@@
5:        ${\bm{v}}_{1}(n)\leftarrow{\bm{I}}_{n}{\mathbf{u}}$
6:        ${\bm{v}}_{2}(n)\leftarrow{\bm{I}}_{n}{\mathbf{v}}$
7:        $d_{1}^{2}(n)\leftarrow\langle{\mathbf{u}},{\bm{v}}_{1}(n)\rangle$
8:        $d_{2}^{2}(n)\leftarrow\langle{\mathbf{v}},{\bm{v}}_{2}(n)\rangle$
@@
16:     $\Delta{\mathbf{u}}\leftarrow\sum_{n=1}^{N}\left[r(n)-\bar{r}\right]\left[{\bm{u}}_{1}(n)-\bar{\bm{u}}_{1}\right]$
17:     $\Delta{\mathbf{v}}\leftarrow-\sum_{n=1}^{N}\left[r(n)-\bar{r}\right]\left[{\bm{u}}_{2}(n)-\bar{\bm{u}}_{2}\right]$
18:     ${\mathbf{u}}\leftarrow{\mathbf{u}}+\eta\Delta{\mathbf{u}}$
19:     ${\mathbf{v}}\leftarrow{\mathbf{v}}+\eta\Delta{\mathbf{v}}$
20:     ${\mathbf{u}}\leftarrow\alpha{\mathbf{u}}/\|{\mathbf{u}}\|$
21:     ${\mathbf{v}}\leftarrow\alpha{\mathbf{v}}/\|{\mathbf{v}}\|$
@@ Appendix C Methods: Early Visual Model Experiments
C.1 Dataset
Distortions were generated for images from the Kodak TID 2008 dataset (Ponomarenko et al., 2009).
@@ Appendix D Methods: Deep Neural Network Experiments
With the exception of the experiment described in the next paragraph, we analyzed DNNs were obtained from the model loading code and checkpoints available at https://github.com/jenellefeather/model_metamers_pytorch which were used in (Feather et al., 2023), and allowed for easy loading and selection of the intermediate layer stages for many models that had previously been proposed as models of human visual perception. The checkpoints for the standard ResNet50 and AlexNet models were obtained from the public PyTorch checkpoints (Ansel et al., 2024). The Stylized Image Net AlexNet (alexnet_trained_on_SIN) and ResNet50 (resnet50_trained_on_SIN) were obtained from https://github.com/rgeirhos/texture-vs-shape associated with (Geirhos et al., 2019). The checkpoint for the ResNet50 $\ell_{2}(\epsilon_{\text{train}}=3.0)$ adversarially trained model was obtained from (Engstrom et al., 2019), and the checkpoint for the Alexnet $\ell_{2}(\epsilon_{\text{train}}=3.0)$ adversarially trained model was obtained from (Feather et al., 2023). For all experiments, we only included intermediate layers before the final classification stage in the principal distortion analysis, and the subset of layers followed those chosen in Feather et al. (2023). Specifically, for the AlexNet models we included layers relu0, relu1, relu2, relu3, relu4, fc0_relu, and fc1_relu in each set of anlayses. For the ResNet50 models we included conv1_relu1, layer1, layer2, layer3, layer4, and avgpool in each set of analyses.
To demonstrate that our approach holds for more modern architectures, we compared a ViT vs. and EfficientNet model. These models were obtained from the PyTorch Image Models (timm) repository (Wightman, 2019). We chose a version of EfficientNet and ViT that were trained on the size of Images in ImageNet-9 ($224\times 224$ images), and versions of each architecture trained with the same image dataset (ImageNet-1k). For the EfficientNet, we tested model EfficientNet-B0 (timm model tf_efficientnet_b0) and included layers conv_stem, blocks.0, blocks.1, blocks.2, blocks.3, blocks.4, blocks.5, blocks.6, and global_pool in the analysis. For the ViT, we tested model ViT (Base-Patch16-224) (timm model vit_base_patch16_224.augreg_in1k) and included layers patch_embed, blocks.0, blocks.1, blocks.2, blocks.3, blocks.4, blocks.5, blocks.6, blocks.7, blocks.8, blocks.9, blocks.10, blocks.11 and head in the analysis.
D.1 Datasets
@@
For the deep neural network experiments we had two sets of images that were used to generate the principal distortions. For the main experiments, we use a subset of 100 ImageNet images where each image was chosen from a unique class (randomly chosen from the set of images at https://github.com/EliSchwartz/imagenet-sample-images).
For the experiments with background and foreground blur, we used the ImageNet-9 dataset with foreground/background masks (Xiao et al., 2021). We chose eight random images from each category of images, resulting in 72 total images. As this is a relatively small subset of images, we removed images that had blank backgrounds (by eye) or that seemed to have incorrect masks (as the masks were automatically generated in the original dataset and not validated with human labeling) and replaced them with new random images so that the backgrounds would always be affected by the smoothing procedure. This procedure was done before running the images through the principal distortion analysis. For the foreground and background blurred images, we blurred the image with a Gaussian filter with a standard deviation of 5 pixels.
D.2 Principal distortion optimization
For each image and each comparison, we ran the gradient descent procedure for principal distortion optimization for 2500 iterations, using an exponentially decaying learning rate that started at 10.0 and decayed to 0.001 by the final step. The exception to this is for the experiment with ViT (Base-Patch16-224) vs. EfficientNet-B0, where we ran the optimization for only 500 iterations.
@@ D.2 Principal distortion optimization
We used a target distortion size of $\alpha=0.1$ and at each step of the optimization, we also scaled ${\bm{\epsilon}}$ so that the image ${\mathbf{s}}+1000{\bm{\epsilon}}$ would not be clipped when the RGB value was represented between 0–1; that is, we scaled ${\bm{\epsilon}}$ so that $0\leq s[i]+1000\epsilon[i]\leq 1$ for each value $i$ in the image. This constraint could potentially bias the perturbations to be more spread out across the image (because all of the amplitude for the perturbation cannot be focused at a small set of pixels). However, removing this constraint did not lead to quantitatively different results (Supp. Fig. SI.10), but the inclusion allows for more viable comparisons with human perception (as the perturbation can be scaled while maintaining valid values in the image gamut). Finally, if one or more of the FIMs is degenerate (i.e., not full rank), this can potentially lead to issues in the optimization procedure (though this issue did not arise in the experiments presented here). To address this, one option is to regularize the FIM by adding a small constant times the identity matrix.
D.3 Supplemental Figures
Figure SI.1: Illustration of method for comparing early visual models to human perception. A) An example base image (${\mathbf{s}}$) and the principal distortions (${\bm{\epsilon}}_{1}$, ${\bm{\epsilon}}_{2}$) generated for the four early visual models (as in Fig. 2). B) Base image corrupted by the two principal distortions (top and bottom row), scaled by amplitudes that sum to 50 and correspond to different log sensitivity ratios (parenthesized values below the pairs, also indicated with grey arrows in panel C). A perceptual experiment can be designed to test which pair of corrupted images (indexed A through I) appear equally distorted to a human observer. This could be done directly, by showing various pairs of distorted images (i.e., ${\mathbf{s}}+k_{1}{\bm{\epsilon}}_{1}$ and ${\mathbf{s}}+k_{2}{\bm{\epsilon}}_{2}$ with varying amplitudes $k_{1},k_{2}$) and asking an observer which of the two appears more distorted with respect to the original ($\mathbf{s}$). The pair of images for which an observer cannot decide (i.e., gives random answers) corresponds to the observer’s log sensitivity ratio. Alternatively, this could be assessed by measuring the observer’s detection thresholds for each distortion independently, and then taking their log ratio. This estimated human log sensitivity ratio can then be compared to the models’ log sensitivity ratios (colored dots, panel C), to assess which model is best aligned with human behavior. Images are best viewed at high resolution.
Figure SI.2: Comparison of ImageNet-1k trained EfficientNet and Vision Transformer. A) Example principal distortions and associated log sensitivity ratio plots when comparing layers of EfficientNet-B0 and ViT (Base-Patch16-224). B) When measured over 72 images from the ImageNet validation set (chosen from ImageNet-9, Xiao et al., 2021), the obtained principal distortions reliably separate the models by architecture—the layers of EfficentNet are more sensitive to distortion ${\bm{\epsilon}}_{1}$, while the layers of ViT are more sensitive to ${\bm{\epsilon}}_{2}$. The principal distortions are qualitatively different than those obtained when comparing AlexNet and ResNet50—for instance, the principal distortions ${\bm{\epsilon}}_{2}$ (i.e., the distortion that the ViT layers are more sensitive) have notable grid artifacts corresponding to the size of the input patches to the ViT model. C) The distortions modify or preserve the label computed by each model, consistent with their predicted sensitivity (see Supp. Fig. SI.4 for a detailed explanation).